If A and B are independent events with ?(?) = 1 4 and ?(?) = 2 3 , calculate the following probabilities:

a. ?(?|?) b. ?(? AND ?)

You roll a dice twice. Let A be the event that you roll a dice and the number shown is 5. Let B be the event that you roll a dice and the number shown is 2. Interpret ?(?|?) in words.

If ?(? ??? ?) = 0.15, and ?(?) = 0.6, what is ?(?|?)?
Answer #1

ANSINER o Criven that - 2 - P(B)= 3 b) P (A AND B) = P(A/B) P(!) + P (B) - P (AUB) Here At B are independent Hen P(A/B) = P(A